# Bidirectional Generative Pre-training for Improving Time Series Representation Learning

### v1

@@ Abstract
Learning time-series representations for discriminative tasks has been a long-standing challenge. Current pre-training methods are limited in either unidirectional next-token prediction or randomly masked token prediction. We propose a novel architecture called Bidirectional Timely Generative Pre-trained Transformer (BiTimelyGPT), which pre-trains on time-series data by both next-token and previous-token predictions in alternating transformer layers. This pre-training task preserves original distribution and data shapes of the time-series. Additionally, the full-rank forward and backward attention matrices exhibit more expressive representation capabilities. Using biosignal data, BiTimelyGPT demonstrates superior performance in predicting neurological functionality, disease diagnosis, and physiological signs. By visualizing the attention heatmap, we observe that the pre-trained BiTimelyGPT can identify discriminative segments from time-series sequences, even more so after fine-tuning on the task.
Machine Learning, ICML
@@ 1 Introduction
Time-series data can provide valuable insights into patient health status and their clinical outcomes (Ma et al., 2023; Eldele et al., 2021). One useful application is to learn temporal representations and use them to achieve accurate sequence-to-vector (seq2vec) prediction tasks. In particular, biosignals such as electroencephalogram and electrocardiogram are extensively utilized for diverse applications such as disease diagnosis (Phan et al., 2021), neurophysiological functionality (Stirling et al., 2020), and vital signal estimation (Reiss et al., 2019b). A long-standing challenge lies in learning meaningful representations from the time-series, particularly in scenarios with limited labeled data. To address this problem, self-supervised learning designs pre-training tasks to learn representations from unlabeled time-series data, thereby improving performance in downstream tasks (Ma et al., 2023).
@@
Figure 1: Three pre-training methods on a normalized time-series with a shapelet learned by pyts package (Faouzi & Janati, 2020). a. Generative pre-training maintains the original distribution and time-series shapelet. b. Masking-based pre-training masks segments to zeros, resulting in distribution shift. c. Dropping-based pre-training randomly discards segments up to 70% of the data and thus disrupts the discriminative shapelet.
@@
Recent advancements in Transformer pre-trained models (PTMs) for Natural Language Processing and Computer Vision have led to their exploration in time-series representation learning. Self-supervised learning methods in this domain fall primarily into two categories: masking-based and dropping-based pre-training. Masking-based methods, such as PatchTST, randomly mask segments of time-series to zeros and then reconstruct them using learned representations (Zerveas et al., 2021; Nie et al., 2023). However, as shown in Figure 1.b, these methods suffer from data distribution shift, which introduce noises into the pre-training stage and degrade the quality of the learned representations. Unlike images and text data (Devlin et al., 2019; Dosovitskiy et al., 2021), distribution shift presents a larger challenge for time-series, where the introduced zeros are substantially deviate from the actual signals (Zhang et al., 2023). In contrast, Cross-Reconstruction Transformer (CRT) adopts a dropping-based method to circumvent distribution shift by discarding and reconstructing data segments (Zhang et al., 2023). However, as shown in Figure 1.c, this method risk disrupting shapelets, which are discriminative subsequences of time-series data crucial for identifying the target class labels of the sequence (Ye & Keogh, 2009; Grabocka et al., 2014). As a potential solution, GPT adopts a next-token prediction task, which effectively preserves both the original distribution and time-series shapelets without changing the data (Figure 1.a). Despite GPT’s efficacy in representation learning in diverse domains, its potential in time-series domain is still underexplored (Radford & Narasimhan, 2018).
@@
While GPT is exceptionally effective for generative forecasting tasks using past contexts, its causal attention (Figure 2.b) is less effective in discriminative tasks that require bidirectional contexts (Devlin et al., 2019). Using bidirectional self-attention (Figure 2.a), PatchTST and CRT achieve superior performance in these tasks. While this self-attention mechanism captures bidirectional contexts, its attention matrix is low rank and thus faces a loss of expressiveness (Bhojanapalli et al., 2020). These limitations underscore the necessity for an architecture that extracts bidirectional contexts with an expressive, full-rank attention matrix.
@@
In this study, we aim to improve time-series representation learning via a novel pre-training strategy towards improving downstream seq2vec prediction tasks. To this end, we propose the Bidirectional Timely Generative Pre-trained Transformer (BiTimelyGPT) that integrates bidirectionality into generative pre-training to improve time-series representation learning. It introduces a Next-Previous-Token Prediction pre-training task, preserving the original distribution and time-series shapelets without any data alterations. Our BiTimelyGPT model alternates forward and backward attention across layers, pre-training deep bidirectional representations through layers. Furthermore, both the forward and backward attention matrices in BiTimelyGPT are full-rank and thus exhibit expressive representation power (Bhojanapalli et al., 2020). The qualitative and quantitative results demonstrate the strengths of BiTimelyGPT over existing studies (Table 1), including data preservation, bidirectionality, and full-rankedness. This makes BiTimelyGPT a highly effective architecture for downstream discriminative tasks.
The key contributions of our research are threefold:
@@
    a Next-Previous-Token Prediction pre-training task preserves the original data distribution and time-series shapelets (Figure 1);
@@
    a Bidirectionally Alternating AutoRegressive Modeling (BAAR) framework alternately models left-to-right and right-to-left information, learning deep bidirectional contexts for discriminative tasks (Figure 2);
@@
    the full-rank forward and backward attention matrices exhibit expressive representation power (Figure 2).
2 Background
@@ 2.1 Transformer Architecture
Transformer is an encoder-decoder model architecture stacked by $L$ layers of Transformer blocks (Vaswani et al., 2017). Each block consists of a multi-head self-attention layer followed by a feed-forward layer. Given an input embedding ${\bm{X}}\in\mathbb{R}^{N\times d}$, where $N$ is the number of tokens and $d$ is the hidden dimension, the output of a self-attention layer is computed as follows.
@@
Among various transformers, GPT and BERT stand out as decoder-only and encoder-only architectures, respectively (Radford et al., 2019; Devlin et al., 2019). The causal Transformer GPT utilizes a masked self-attention mechanism:
@@
where both the mask matrix ${\bm{C}}$ and forward attention matrix $\overrightarrow{{\bm{A}}}$ are lower triangular. GPT’s Next-Token Prediction task effectively captures complex dependencies of natural language and thus excels in generative tasks. In contrast, BERT uses masking-based pre-training for deep bidirectional representations, which are crucial for discriminative tasks. Unlike BERT’s bidirectional attention in each layer, BiTimelyGPT alternates between forward and backward attention across layers, pre-training deep bidirectional representations through a generative task.
Figure 3: BiTimelyGPT overview. a. BiTimelyGPT architecture is stacked by a series of bidirectional retention layers, with an overflow of pre-training in Section A.2. b. BAAR framework (Section 3.1) alternately models left-to-right and right-to-left information across layers. It performs Next-Previous-Token Prediction on the final two layers and utilizes the [SOS] token in the final layer for discriminative tasks. c. Bidirectional Retention Layer (Section 3.2) alternates between the forward and backward retention mechanism across layers.
@@ 2.2 Bidirectional Recurrent Representation Learning
Deep Bidirectional LSTM (BiLSTM) trains forward and backward LSTMs independently at each layer, followed by concatenating the hidden states from both directions post-layer. Embeddings from Language Models (ELMo) also trains forward and backward LSTMs independently, but concatenates their outputs only in the final layer (Peters et al., 2018). Although ELMo effectively generates pre-trained contextualized representations for downstream tasks, its bidirectional context is not deep due to the shallow concatenation of left and right contexts (Devlin et al., 2019). In contrast, our proposed BAAR framework alternately models information from both directions across layers, allowing for deep bidirectional representation learning.
@@ 2.3 Recurrent Transformer RetNet
Extrapolatable position embedding (xPos) encodes relative position information into Query $\tilde{\bm{Q}}_{n}$ and Key $\tilde{\bm{K}}_{m}$ based on the distances between tokens $n$ and $m$ (Sun et al., 2023b):
@@
where $\theta$ and $\gamma$ are rotation and exponential decay hyperparameters. The RetNet model introduces a retention mechanism based on the xPos embedding: (Sun et al., 2023a):
@@
where $\overrightarrow{{\bm{D}}}\in\mathbb{R}^{N\times N}$ is a lower triangular decay matrix. The retention can be reformulated as an RNN for $n$-th timestep:
@@
TimelyGPT, built upon RetNet, offers two main advantages in ultra-long-term forecasting (Song et al., 2023). Firstly, TimelyGPT leverages xPos to encode trend and periodic temporal patterns into time-series representations. Moreover, its retention mechanism, reformulating itself as an RNN, inherently captures temporal dependencies. However, unidirectional TimelyGPT falls short in discriminative tasks compared to bidirectional CRT without convolution modules. To bridge this gap, we propose BiTimelyGPT, which leverages a novel BAAR framework, to integrate bidirectionality into autoregressive time-series modeling.
@@ 3 BiTimelyGPT Methodology
An input sequence of time-series $x$ is denoted as $\{x_{1},x_{2},\ldots,x_{T}\}$ over $T$ timesteps, with $x_{t}$ comprising $V$ features. As depicted in Figure 3.a, BiTimelyGPT utilizes a convolution-subsampling tokenizer with two 1-D convolution layers (Gulati et al., 2020), reducing the $T\times V$ down to an $N\times V$ sequence with $N<T$. These tokens are projected onto an input embedding ${\bm{X}}\in\mathbb{R}^{N\times d}$, which is prepended and appended with [SOS] and [EOS] tokens (Radford et al., 2021), respectively. Section A.2 depicts the details. As stated above, the goal of time-series representation learning in this study is to pre-train temporal representations for fine-tuning seq2vec tasks. Here, the output vector $\hat{y}$ can be either a continuous variable for regression or a categorical variable for classification. We describe the BAAR framework in Section 3.1, the BiTimelyGPT pre-training in Section 3.2, and the desired full-ranked attention matrix as a result in Section 3.3. The notations of variables are defined in Section A.1.
@@ 3.1 Bidirectional Alternating AutoRegressive Modeling
The BAAR framework presents a general approach for incorporating bidirectionality into autoregressive models. It alternates between left-to-right and right-to-left information flows depending on the odd and even layer indices, respectively. In the odd-indexed layers $l\in\{1,3,\ldots,L-1\}$, the forward model computes the probability of the sequence $x$ by modeling the conditional probability of $x_{t}$ given the history context $(x_{1},\ldots,x_{t-1})$ (Bengio et al., 2000):
@@
This results in a history-dependent representation vector $\overrightarrow{h}_{t}^{(l)}$ given its history context $(x_{1},\ldots,x_{t-1})$. In the even-indexed layers $l\in\{2,4,\ldots,L\}$, the backward model runs over the sequence $x$ in reverse, predicting a previous timestep $x_{t}$ given the future context $(x_{t+1},\ldots,x_{T})$:
@@
This yields a future-dependent representation vector $\overleftarrow{h}_{t}^{(l)}$ given its future context $(x_{t+1},\ldots,x_{T})$. The conditional probabilities in Equation 6 and Equation 7 can be modeled by forward and backward attention (Radford & Narasimhan, 2018).
Figure 3.b illustrates how the BAAR framework alternates between forward and backward layers to learn deep bidirectional representations. In the odd-indexed layer, the forward attention starts with a [SOS] token to predict the next tokens from left to right, where the [EOS] token summarizes information from all timesteps. In the even-indexed layer, the backward attention begins with the [EOS] token to predict the previous token in the right-to-left direction, intending to reconstruct the sequence in reverse. By alternating left-to-right and right-to-left information flows, the [EOS] and [SOS] tokens aggregate the sequence’s features in the odd and even layers, respectively. In the final layer (assuming $L$ is an even number), the [SOS] token serves as a sequence representation for discriminative tasks.

### v3

# Bidirectional Generative Pre-training for Improving Healthcare Time-series Representation Learning

@@ Abstract
Learning time-series representations for discriminative tasks, such as classification and regression, has been a long-standing challenge in the healthcare domain. Current pre-training methods are limited in either unidirectional next-token prediction or randomly masked token prediction. We propose a novel architecture called Bidirectional Timely Generative Pre-trained Transformer (BiTimelyGPT) [^1] We have open-sourced source code at https://github.com/li-lab-mcgill/BiTimelyGPT, which pre-trains on biosignals and longitudinal clinical records by both next-token and previous-token prediction in alternating transformer layers. This pre-training task preserves original distribution and data shapes of the time-series. Additionally, the full-rank forward and backward attention matrices exhibit more expressive representation capabilities. Using biosignals and longitudinal clinical records, BiTimelyGPT demonstrates superior performance in predicting neurological functionality, disease diagnosis, and physiological signs. By visualizing the attention heatmap, we observe that the pre-trained BiTimelyGPT can identify discriminative segments from biosignal time-series sequences, even more so after fine-tuning on the task.
@@ 1 Introduction
Healthcare time-series data can provide valuable insights into patient health status and their clinical outcomes (Ma et al., 2023; Eldele et al., 2021). In the field of healthcare, there are two primary categories: continuous and irregularly sampled time-series data. Continuous time-series, such as biosignals, has been extensively utilized for diverse applications such as disease diagnosis (Phan et al., 2021), neurophysiological functionality (Stirling et al., 2020), and vital signal estimation (Reiss et al., 2019). Irregularly sampled time series is commonly found in electronic health records (EHRs), where spontaneous updates are made due to outpatient visits or inpatient stays (Zhang et al., 2022a). One valuable application is learning temporal representations from healthcare data to enhance sequence-to-vector (seq2vec) prediction tasks. A long-standing challenge lies in learning meaningful representations from the time-series, particularly in scenarios with limited labeled data. To address this problem, self-supervised learning designs pre-training tasks to learn representations from unlabeled time-series data, thereby improving performance in downstream tasks (Ma et al., 2023).
@@
Figure 1: Three pre-training methods on a normalized biosignal sequence. The green segment denotes a discriminative subsequence (shapelet) learned by a published pyts package (Faouzi and Janati, 2020). a. Generative pre-training preserves its original distribution and time-series shapelet, maintaining the integrity of the sequence information. b. Masking-based pre-training masks segments as zeros, resulting in an issue of distribution shift. c. Dropping-based pre-training randomly discards up to 70% of the segments and thus disrupts its discriminative shapelet.
@@
Recent advancements in Transformer pre-trained models (PTMs) for Natural Language Processing and Computer Vision have led to their exploration in time-series representation learning. Self-supervised learning methods in this domain fall primarily into two categories: masking-based and dropping-based pre-training. Masking-based methods, such as PatchTST, randomly mask segments of time-series as zeros and then reconstruct them using learned representations (Zerveas et al., 2021; Nie et al., 2023). However, as shown in Figure 1.b, these methods suffer from data distribution shift, which introduces noise into the pre-training stage and degrades the quality of the learned representations. Unlike images and text data (Devlin et al., 2019; Dosovitskiy et al., 2021), distribution shift presents a larger challenge for time-series data, where the introduced zeros substantially deviate from the actual signals (Zhang et al., 2023). In contrast, Cross-Reconstruction Transformer (CRT) adopts a dropping-based method to circumvent distribution shift by discarding and reconstructing data segments (Zhang et al., 2023). However, as shown in Figure 1.c, this method risks disrupting shapelets, which are discriminative subsequences of time-series data crucial for identifying the target class labels of the sequence (Ye and Keogh, 2009; Grabocka et al., 2014). As a potential solution, GPT adopts a next-token prediction task, which effectively preserves both the original distribution and time-series shapelets without altering the data (Figure 1.a). Despite GPT’s efficacy in representation learning in diverse domains, its potential in time-series domain is still underexplored (Radford and Narasimhan, 2018).
@@
While GPT is exceptionally effective for generative forecasting tasks using the unidirectional history contexts, its causal attention (Figure 2.b) is less effective in discriminative tasks that require bidirectional contexts (Devlin et al., 2019). Using bidirectional self-attention mechanism (Figure 2.a), PatchTST and CRT achieve superior performance in these discriminative tasks. While this self-attention mechanism captures bidirectional contexts, its attention matrix is low rank and thus faces a loss of expressiveness (Bhojanapalli et al., 2020). These limitations underscore the necessity for a novel architecture that extracts bidirectional contexts with an expressive, full-rank attention matrix.
@@
In this study, we aim to improve healthcare time-series representation learning via a novel pre-training strategy towards enhancing downstream seq2vec prediction tasks. To this end, we propose the Bidirectional Timely Generative Pre-trained Transformer (BiTimelyGPT) that integrates bidirectionality into generative pre-training to improve time-series representation learning. It introduces a Next-Previous-Token Prediction pre-training task, preserving the original distribution and time-series shapelets without any data alterations (Figure 1.a). BiTimelyGPT alternates forward and backward attention across layers to pre-train deep bidirectional representations. Furthermore, both forward and backward attention matrices in BiTimelyGPT are full-rank and thus exhibit expressive representation power (Bhojanapalli et al., 2020). The qualitative and quantitative results demonstrate the strengths of BiTimelyGPT over existing methods (Table 1), including data preservation, bidirectionality, and full-rankedness. The key contributions are threefold:
@@
    a novel Next-Previous-Token Prediction pre-training task preserves the original data distribution and time-series shapelets (Figure 1);
@@
    a novel Bidirectionally Alternating AutoRegressive Modeling (BAAR) framework alternately models left-to-right and right-to-left information across layers, learning deep bidirectional contexts for discriminative tasks (Figure 3);
@@
    the resulting forward and backward attention matrices, which are full-rank, exhibit expressive representation capabilities for pre-training (Figure 2).
Generalizable insights about machine learning in the context of healthcare
Healthcare time series provide rich opportunities for learning representations that enhance downstream medical applications, such as disease classification and physical activity prediction. Our study proposes BiTimelyGPT to effectively pre-train temporal representations from biosignals and longitudinal clinical records and provides two key insights into modeling time-series health data:
  1. 1.
    We have demonstrated the significance of modeling bidirectional contexts with full-rank attention matrices for predicting both continuously monitored physiological measurements and irregularly sampled diagnostic records. Our finding reveals that representation learning of complex temporal patterns can enhance downstream discriminative tasks. Future research could explore the generalizability of representation learning across different types of healthcare data and various downstream tasks.
  2. 2.
    We have developed a flexible and efficient algorithm, BiTimelyGPT, for representing deep bidirectional contexts in various healthcare time series data. This approach outperforms current state-of-the-art techniques in modeling and predicting physiological signals and longitudinal medical records. Future research can make use of this algorithm or design novel approaches based on our study to further enhance discriminative tasks in different healthcare datasets.
2 Related Work
@@ 2.1 Transformer architecture
Transformer is an encoder-decoder model architecture stacked by $L$ layers of Transformer blocks (Vaswani et al., 2017). Each block consists of a self-attention layer followed by a feed-forward layer. Given an input embedding ${\bm{X}}\in\mathbb{R}^{N\times d}$, where $N$ is the number of tokens and $d$ is the hidden dimension, the output of a self-attention layer is computed as follows.
@@
Among various Transformer models, GPT and BERT stand out as decoder-only and encoder-only architectures, respectively (Radford et al., 2019; Devlin et al., 2019). The unidirectional GPT utilizes a causal self-attention mechanism:
@@
where both the mask matrix ${\bm{C}}$ and forward attention matrix $\overrightarrow{{\bm{A}}}$ are lower triangular (Figure 2.b). GPT’s Next-Token Prediction task effectively captures complex dependencies of natural language and thus excels in generative tasks. In contrast, BERT uses the self-attention (Figure 2.a) and masking-based pre-training for deep bidirectional representations, which are crucial for discriminative tasks. Unlike BERT’s bidirectional attention in each layer, BiTimelyGPT alternates between forward and backward attention across layers (Figure 2.b,c), learning deep bidirectional representations via a generative pre-training.
@@ 2.2 Bidirectional recurrent representation learning
Deep Bidirectional Long Short-Term Memory (BiLSTM) independently trains forward and backward LSTMs at each layer, followed by concatenating the hidden states from both directions post-layer. Embeddings from Language Models (ELMo) also trains forward and backward LSTMs independently, but concatenates their outputs only in the final layer (Peters et al., 2018). Although ELMo effectively extracts pre-trained contextualized representations for downstream tasks, its bidirectional context is not deep due to the shallow concatenation of left and right contexts (Devlin et al., 2019). In contrast, we propose a BAAR framework that alternately models information flows from both directions across layers, allowing for deep bidirectional representation learning.
@@ 2.3 Recurrent Transformer RetNet
Extrapolatable position (xPos) embedding encodes relative position information into Query $\tilde{\bm{Q}}_{n}$ and Key $\tilde{\bm{K}}_{m}$ based on the distances between tokens $n$ and $m$ (Sun et al., 2023b):
@@
where $\theta$ and $\gamma$ are rotation and exponential decay hyperparameters. The RetNet model introduces a Retention mechanism based on the xPos embedding (Sun et al., 2023a):
@@
where $\overrightarrow{{\bm{D}}}\in\mathbb{R}^{N\times N}$ is a lower triangular decay matrix. The Retention mechanism can be reformulated as an RNN for $n$-th timestep:
@@
TimelyGPT, built upon RetNet, performs long-term forecasting for biosignals and longitudinal clinical records (Song et al., 2023). To adapt irregularity in EHR data, TimelyGPT adapts the relative distance $n–m$ to account for the uneven gap between diagnoses, enabling it to compute time-decaying attention for these irregular gaps. Consequently, each element of the decay matrix $\overrightarrow{D}_{nm}$ in Equation 4 reflects the relative distance between the irregularly sampled timesteps $n$ and $m$ in months. However, unidirectional TimelyGPT falls short in discriminative tasks compared to bidirectional CRT without the aid of convolution modules. To bridge this gap, we propose BiTimelyGPT, which leverages a novel BAAR framework, to integrate bidirectionality into autoregressive time-series modeling.
@@ 3 BiTimelyGPT Methodology
An input sequence of time-series $x$ is denoted as $\{x_{1},x_{2},\ldots,x_{T}\}$ over $T$ timesteps, with $x_{t}$ comprising $V$ features. As depicted in Figure 3.a, BiTimelyGPT utilizes a convolution-subsampling tokenizer with two 1-D convolution layers, reducing the $T\times V$ down to an $N\times V$ sequence with $N<T$. These tokens are projected onto an input embedding ${\bm{X}}\in\mathbb{R}^{N\times d}$, which is prepended and appended with [SOS] and [EOS] tokens, respectively. Section A.2 depicts the details. As stated above, the goal of time-series representation learning in this study is to pre-train temporal representations for fine-tuning seq2vec prediction tasks. Here, the output vector $\hat{y}$ can be either a continuous variable for regression or a categorical variable for classification. We describe the BAAR framework in Section 3.1, the BiTimelyGPT pre-training in Section 3.2, and the desired full-rank attention matrix as a result in Section 3.3. The notations of variables are defined in Section A.1.
Figure 3: BiTimelyGPT overview. a. BiTimelyGPT architecture is stacked with $L$ bidirectional Retention layers, with an overflow of pre-training in Section A.2. b. BAAR framework (Section 3.1) alternately models left-to-right and right-to-left information across layers. It performs Next-Previous-Token Prediction on the final two layers and utilizes the [SOS] token in the final layer for discriminative tasks. c. Bidirectional Retention Layer (Section 3.2) alternates between forward and backward Retention across layers.
@@ 3.1 Bidirectional Alternating AutoRegressive Modeling
The BAAR framework presents a general approach for incorporating bidirectionality into autoregressive models. It alternates between left-to-right and right-to-left information flows depending on the odd and even layer indices, respectively. In the odd-indexed layers $l\in\{1,3,\ldots,L-1\}$, the forward model computes the probability of the sequence $x$ via a conditional probability of $x_{t}$ given the history context $(x_{1},\ldots,x_{t-1})$ (Bengio et al., 2000):
@@
This results in a history-dependent representation vector $\overrightarrow{h}_{t}^{(l)}$ given its history context $(x_{1},\ldots,x_{t-1})$. In the even-indexed layers $l\in\{2,4,\ldots,L\}$, the backward model runs over the sequence $x$ in reverse, predicting a timestep $x_{t}$ given the future context $(x_{t+1},\ldots,x_{T})$:
@@
This yields a future-dependent representation vector $\overleftarrow{h}_{t}^{(l)}$ given its future context $(x_{t+1},\ldots,x_{T})$. The conditional probabilities in Equation 6 and Equation 7 can be modeled by forward and backward attention (Radford and Narasimhan, 2018).
Figure 3.b illustrates how the BAAR framework alternates between forward and backward layers to learn deep bidirectional representations. In the odd-indexed layer, the forward attention starts with a [SOS] token to predict the next tokens from left to right, where the [EOS] token summarizes information from all timesteps. In the even-indexed layer, the backward attention begins with the [EOS] token to predict the previous token in the right-to-left direction, intending to reconstruct the sequence in reverse. By alternating left-to-right and right-to-left information flows, the [EOS] and [SOS] tokens aggregate the sequence’s features in the odd and even layers, respectively. In final layer (assuming an even number $L$), the [SOS] token serves as a sequence representation for discriminative tasks.